OpenStudy (anonymous):

log(5)x+log(5)(x+3)=1 log(5)x^2+3x=1 5^1=x^2+3x x^2+3x-5=0 (x+?)(x-?)=0

OpenStudy (anonymous):

Well using the quadratic formula,you could factor \[x^2+3x-5=0\] to get, \[ \left\{\left\{x=\frac{1}{2} \left(-3-\sqrt{29}\right)\right\},\left\{x=\frac{1}{2} \left(-3+\sqrt{29}\right)\right\}\right\} \]

OpenStudy (anonymous):

I dont think thats correct...?

OpenStudy (anonymous):

log(5)x + log(5)(x+3) = 1 log(5) (x^2 + 3x) = 1 x^2 + 3x = 5 x^2 + 3x - 5 = 0 x = (-3 +/- sqrt(9 + 20)) / 2 foolformath is right..